OpenStudy (anonymous):

When graphing y < 5x - 2, what kind of boundary line is used and where will the graph be shaded?

OpenStudy (jdoe0001):

you'd end up graphing the line y = 5x-2 the line is DASHED when using > or < the line is SOLID when using \(\bf \ge \ or \ \le\)

jimthompson5910 (jim_thompson5910):

also, you a) shade above the boundary line if you have > or \(\large \ge\) OR b) shade below the boundary line if you have < or \(\large \le\)